OpenStudy (anonymous):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost? $75 $100 $125 $150

OpenStudy (anonymous):

Cost of Hotel = H Cost of Rent = R Days Rented = D Nights Rented= N How much Paid = P So.... P = H x N + R x D Connor's: $675 = H x 3 nights + R x 4 days Jillian's: $875 = H x 4 nights + R x 5 days Solve: ($675 = H x 3 nights + R x 4 days) x (5) = $3,375 = 15 nights x H + 20 days x R ($875 = H x 4 nights + R x 5 days) x (-4) = - $3,500 = -16 nights x H + -20 days x R $3,375 = 15 nights x H + 20 days x R - $3,500 = -16 nights x H + -20 days x R -$125 = -1 nights x H H = $125

OpenStudy (anonymous):

;)